### v1

# PADE: A Predictor-Free Sparse Attention Accelerator via Unified Execution and Stage Fusion

Abstract
Attention-based models have revolutionized AI, but the quadratic cost of self-attention incurs severe computational and memory overhead. Sparse attention methods alleviate this by skipping low-relevance token pairs. However, current approaches lack practicality due to the heavy expense of added sparsity predictor, which severely drops their hardware efficiency.
This paper advances the state-of-the-art (SOTA) by proposing a bit-serial enable stage-fusion (BSF) mechanism, which eliminates the need for a separate predictor. However, it faces key challenges: 1) Inaccurate bit-sliced sparsity speculation leads to incorrect pruning; 2) Hardware under-utilization due to fine-grained and imbalanced bit-level workloads. 3) Tiling difficulty caused by the row-wise dependency in sparsity pruning criteria.
We propose PADE, a predictor-free algorithm-hardware co-design for dynamic sparse attention acceleration. PADE features three key innovations: 1) Bit-wise uncertainty interval-enabled guard filtering (BUI-GF) strategy to accurately identify trivial tokens during each bit round; 2) Bidirectional sparsity-based out-of-order execution (BS-OOE) to improve hardware utilization; 3) Interleaving-based sparsity-tiled attention (ISTA) to reduce both I/O and computational complexity. These techniques, combined with custom accelerator designs, enable practical sparsity acceleration without relying on an added sparsity predictor. Extensive experiments on 22 benchmarks show that PADE achieves $7.43\times$ speed up and $31.1\times$ higher energy efficiency than Nvidia H100 GPU. Compared to SOTA accelerators, PADE achieves $5.1\times$, $4.3\times$ and $3.4\times$ energy saving than Sanger, DOTA and SOFA.
I Introduction
Transformer models have achieved significant success in various fields, spanning from content generation [song2018situ, brown2020language, hendrycks2020measuring, lin2023videodirectorgpt] to computer vision [dosovitskiy2020image, you2023vitcod, bai2024seed]. However, the self-attention mechanism used in Transformers suffers from quadratic time and memory complexity, limiting their scalability to long sequences.
To address this, sparse attention techniques have emerged as a promising solution, where attention is computed over a subset of query-key (Q-K) pairs instead of a dense attention matrix. Existing sparse attention works can be divided into two routes: static sparsity (SS) and dynamic sparsity (DS). SS [parmar2018image, child2019generating, qiu2019blockwise, zaheer2020big, beltagy2020longformer, jiang2024minference, dong2022cswin, katharopoulos2020transformers] relies on predefined sparse patterns that remain fixed during inference, lacking flexibility and often resulting in significant accuracy degradation [li2024large, fuad2023survey, kachris2025survey]. In contrast, DS [correia2019adaptively, kitaev2020reformer, you2023vitcod, wang2023cta, song2024tsacc, cho2024sparc, park2024token, fan2022adaptable, dass2023vitality, li2020ftrans, wang2024sofa, qin2023fact, ham20203, ham2021elsa, lu2021sanger, qu2022dota, yang2022dtatrans, zhou2022energon, wang2025bitstopper, wang2021spatten, liu2022dynamic, wang2025lapa, li2022accelerating, wang2025mcbp] adapts the sparsity pattern at runtime, offering improved accuracy and flexibility, making it more suitable for a broader range of tasks and input types.
Figure 1: Comparison of (a) current DS works and (b) PADE.
[…]
Unfortunately, such an added predictor occupies substantial overhead, increasingly offsetting the sparsity benefit. Fig. 2 (a) shows the power breakdown of dense attention and two representative DS accelerators: Sanger [lu2021sanger], SOFA [wang2024sofa], with varying executor bit-widths. Power consumption is categorized into executor and predictor components. Sanger uses 4-bit MSB multiplication with threshold comparison, while SOFA employs log domain shifting with top-$k$ sorting. As depicted in Fig. 2(a), there are two key observations: (1) At larger executor bit-widths (e.g., 16bit), the DS reduces overall power by about $63\%$, with the predictor costing only $33\%$. This explains why previous DS works adopt additional sparsity predictors. (2) However, as executor bit-width decreases, predictor overhead becomes dominant. At 8-bit, overall savings drop to merely $32\%$, with the sparsity predictor occupying over $63\%$ of total cost. This is due to the predictor must access and process full-sized K tensors, a cost unaffected by sparsity.
Further, Fig. 2(b) reveals the predictor-to-executor power ratio under varying sequence lengths (SL). As can be seen, as the SL increases, this ratio grows noticeably for both designs, indicating the growing relative overhead of the predictor. This is because the increased sparsity in longer sequences exacerbates the predictor’s relative overhead.
Figure 2: (a) Power breakdown of dense and DS attention (SA: Sanger, SO: SOFA) with TSMC 28nm across executor bit-widths of Llama7B. (b) Power ratio of predictor and executor versus SL with under 8-bit quantized executor.
Takeaway: With the rapid advancement of Transformer quantization techniques, like GPTQ [frantar2022gptq], LLM.int8() [dettmers2022gpt3], SmoothQuant [xiao2023smoothquant] and Atom [zhao2024atom], there is a growing trend of adopting low-bit quantization in attention mechanisms. In these cases, the predictor’s overhead increasingly offsets the benefits of sparsity. This highlights a need to reduce or even eliminate the prediction overhead.
Insights: The root cause of the excessive prediction cost stems from the decoupling between existing sparsity predictors and executors, which hinders the computational and memory access efforts paid in the predictor from being reused by the executor. To address this, we draw inspiration from bit-serial computing [judd2016stripes, albericio2017bit, lee2018unpu, chen2024bbs, kam2025panacea, im2023sibia, guo2025transitive], which separates an INT operation into multi-round bit-level steps. This motivates a unified design that integrates prediction and execution into a single computation stage, thereby eliminating the separate predictor and improving overall efficiency.
To realize this idea, we propose a bit-serial-enable stage fusion (BSF) strategy that eliminates the additional prediction stage via the following key steps: 1) Start with the first bit plane (i.e., MSB) of Keys for bit-serial speculating of $\mathbf{Q}\times\mathbf{K}^{T}$. 2) Once a token (Key) is identified as unlikely to be an important QK pair (i.e., iQK), its processing and associated memory access with subsequent bit planes are immediately terminated. In this way, the accelerator only needs to perform the remaining computation for the iQK, and obtains the final result by adding it to the previously generated partial result.
Despite its potential, realizing a BSF-style DS accelerator presents several challenges: (1) Lack of an effective bit-wise decision mechanism for early iEQKs identification. (2) Hardware under-utilization due to fine-grained and imbalanced bit-level workloads. (3) Tiling difficulty arising from row-wise dependency in sparsity pruning criteria.
To this end, we propose PADE, a software-hardware co-design, whose high-level overview is depicted in Fig.1 (b). It features three key techniques that correlate to three challenges:
1) We propose Bit-level Uncertainty Interval-enabled Guarded Filtering (BUI-GF) to accurately identify iEQKs at each bit round. By exploiting properties of two’s complement representation, we define a Bit-wise Uncertainty Interval (BUI) to bound the potential variation of inner products. This safety margin enables precise and reliable early pruning decisions.
Figure 3: Illustration of the DS attention mechanism.
Figure 4: (a) Traditional DS works, featuring stage splitting. (b) Our work features stage-fusion. (c) Reduced complexity for stage splitting and stage fusion.
2) We propose a Bidirectional Sparsity-based Out-of-order Execution (BS-OOE) to improve bit-grained hardware utilization. It first introduces a bit-level bidirectional sparsity (BS) scheme to promote load balancing across PEs. Further, it utilizes out-of-order execution to hide DRAM access latency by breaking the constraint of conventional bit-serial computation.
3) We propose an Interleaving-based Sparsity-Tiled Attention (ISTA) mechanism to improve IO efficiency. By exploiting the monotonicity of Softmax and the early termination property of bit-serial computation, we skillfully decompose pruning decisions to the tiling level. Additionally, an interleaved update strategy further minimizes redundant operations across tiles.
To support the above optimization mechanisms effectively, we design a dedicated accelerator named PADE: 1) For BUI-GF, it employs the dedicated scoreboard-based, result-reusable PE lane to eliminate redundant memory access across bit execution rounds, thus significantly reducing the energy overhead of repeatedly loading bit planes. 2) For BS-OOE, PADE integrates the grouped, lightweight sparsity ANDer trees to mitigate the overhead of large multiplexers. 3) For ISTA, a reuse-aware reorder scheduler is dedicated to improving tiling execution efficiency, by minimizing redundant memory access. The PADE accelerator achieves an average energy efficiency of 11740 GOPS/W, which is 31.1×, 5.1×, 4.3× and 3.4× higher than H100 GPU, SOTA accelerator Sanger, DOTA and SOFA.
II Background
II-A Transformer and DS Attention
Transformer models. Initially, the Transformer maps a length- $S$ sequence into Q, K, and V spaces. Next, Q and K are multiplied to generate an attention score $\mathbf{S}$ with $\mathbb{R}^{S\times S}$, which captures token-to-token correlations. The $\mathbf{S}$ is then passed through a softmax and multiplied with V activation, resulting in a matrix $\mathbf{O}\in\mathbb{R}^{S\times H}$, where $H$ denotes hidden dimension. Finally, a feed-forward network (FFN) generates the outputs.
Dynamic Sparsity (DS) attention. Typically, dense attention involves load and compute all Ks and Vs. In contrast, the DS attention fetches and processes only the important iQK pairs. As depicted in Fig. 3, only the 0th and 3rd KVs are loaded and computed with the Query, thus effectively reducing the computation and memory access. However, this reduced complexity comes at the cost of an additional sparsity predictor, which involves low-bit QK speculation and vital token selection processes. This introduces non-negligible overhead.
III Motivation
III-A Re-examinating DS Works and Opportunity
As shown in Fig. 4(a), traditional DS works adopt a stage-splitting paradigm that decouples prediction from execution. Taking Sanger [lu2021sanger] as an example, the predictor takes the full 4-bit Key tensor to identify iQKs, while the executor separately fetches the corresponding KVs for precise computation. However, this paradigm introduces two inefficiency sources: (1) For Keys related to iEQKs (e.g., K0), a single bit may suffice to identify their insignificance. However, the stage-splitting approach blindly loads 4 bits, resulting in redundancy. (2) For Keys related to iQKs (e.g., Ks), the executor reloads their high-bit-width versions for subsequent more precise computation. However, it fails to reuse the data already processed during prediction, leading to inefficiency.
Quantifying predictor overhead. As shown in Fig. 2, after sparsification, the added predictor incurs over 63% power overhead, limiting the gain of DS attention to just $1.5\times$ over dense attention. To this end, we derive the design guidance:
(Design Guidance) An ideal DS accelerator should eliminate the extra sparsity predictor while preserving sparsity.
Motivated by this insight, the BSF strategy aims to unify prediction and execution within a single computation stage, thereby minimizing both computation and memory access. As illustrated in Fig. 4(b), BSF progressively extracts lower-order bit-planes to assess token importance during QK computation. This enables not only early termination for iEQKs but also computation and memory access reuse for iQKs.
Unlike stage-splitting DS designs that rely on separate predictors, BSF eliminates prediction overhead and enables fine-grained early termination to reduce unnecessary computation and memory access. Fig. 4(c) profiles the memory access and computation reduction achieved by different strategies in the attention modules across four layers from LLaMA-2 7B [touvron2023llama2]. On average, BSF can achieve $4.6\times$ higher memory access and $2.1\times$ more computation reduction, compared to traditional stage-splitting DS approaches, highlighting its significant energy efficiency advantage over current DS methods.
Figure 5: Challenges for bit-serial enable stage fusion. (a)-(b) Inaccuracy (c)-(d) Hardware under-utilization. (e)-(f) Tiling difficulty.
III-B Challenges for Bit-serial Enable Stage-fusion
Despite its theoretical benefits, a naive implementation of BSF will encounter three challenges, as depicted in Fig. 5.
(Challenge 1) Incorrect pruning decisions caused by the inherent inaccuracy in bit-wise speculation.
As shown in the Fig. 5 (a), the 1-bit MSB representation of (+5) and 1-bit representation of (-5) are used to predict the result of (+5) $\times$ (+5) + (+5) $\times$ (-5). The true result should be 0. However, under the 1-bit representation, the MSB bit plane of (1011)2 (-5) and (0101)2 (+5) are regarded as (1000)2 (-8) and (0000)2 (+0), respectively. This leads to an estimated result of -40, which significantly deviates from the correct result. This severe error can lead to incorrect token pruning during early termination. For example, in K7 of Fig. 5 (b).
Key idea We introduce the concept of a bit uncertainty interval (BUI), which defines the maximum fluctuation range of dot products across diverse bit-planes, via simple yet efficient bit-flipping. To support this, we design a scoreboard-based result-reuse PE to reduce BUI hardware overhead and redundant bit-plane memory accesses.
(Challenge 2) Compute resource underutilization due to workload imbalance and exposed memory access latency.
As depicted in Fig. 5 (c), each bit-plane of different Keys is assigned to a separate PE for parallel execution. However, PEs corresponding to Keys whose bit-planes contain more ‘1’ bits will require more processing cycles (e.g., PE 0), while those with fewer ‘1’ bits will finish earlier (e.g., PE 1). This imbalance leads to computation stalls and under-utilization.
As shown in Fig. 5 (d), to enable bit-grained early termination, it is critical to avoid the bulk loading of all bit-planes for each Key. Instead, each Key should be evaluated bit-plane by bit-plane to determine whether it qualifies as an iEQK. If true, the subsequent bit-planes for that Key are skipped; otherwise, the next bit-plane is required. However, due to DRAM’s dynamic precharge mechanism, loading each bit-plane typically incurs several dozen cycles [chang2016understanding, JEDEC_JESD235D]. Naively stalling computation during data loading results in underutilized computational resources.
Key idea. We propose bidirectional sparsity (BS), treating bit ‘1’ as a form of sparsity, thus ensuring load imbalance is below $50\%$. Building upon this, we introduce bit-wise out-of-order (OOE) execution, allowing the PE to process other bit-planes while avoiding memory access stalls. To support this, we propose a lightweight ANDer tree to reduce multiplexer’s overhead from BS sparsity, while the scoreboard-based PE facilitates partial sum buffering and reuse.
(Challenge 3) IO inefficiency resulting from the conflict between tiling with the row-dependent pruning strategy.
As shown in Fig. 5 (e), existing pruning strategies rely on row-wise attention score distributions to assess token importance, introducing strong row dependencies. This dependency prevents effective tiling, which is essential for IO efficiency. Without tiling, when the number of parallel queries increases, memory access overhead grows sharply. As illustrated in Fig. 5 (f), increasing the number of parallel queries ($P$) from 8 to 32, leads to over 12× more memory accesses. A coarse solution is to enlarge on-chip SRAM, but this incurs significant area inefficiency. For example, with (PP=512, SS=2048), 5MB of SRAM is required, resulting in a 5.47 mm2 footprint under TSMC 28nm technology, which is $7.4\times$ and $8.9\times$ larger than the total area of SpAtten [wang2021spatten] and ELSA [ham2021elsa], respectively.
TABLE I: Summary for SOTA Attention Accelerators.
Accelerator Optimization Predictor Tiling Optimiz.
Computation Memory Free Support Level
ELSA [ham2021elsa] $\times$ $\times$ $\times$ Value
Sanger [lu2021sanger] $\times$ $\times$ $\times$ Value
DOTA [qu2022dota] $\times$ $\times$ $\times$ Value
DTATrans[yang2022dtatrans] Low ✓* $\times$ Value
SpAtten [wang2021spatten] Low ✓* $\times$ Multi-bit
Energon [zhou2022energon] $\times$ $\times$ $\times$ Multi-bit
FACT [qin2023fact] $\times$ $\times$ $\times$ Value
SOFA [wang2024sofa] Low $\times$ Value
PADE Bit
  • *  Sparsity guided by preceding layer scores; Accuracy degradation w/o retrain.
Key idea. We reveal and leverage the monotonicity of the softmax function, and further adjust the pruning decision mechanism as follows: retaining tokens that reach the least significant bit (LSB) plane but without being pruned. This enables efficient and I/O-friendly pruning within tiled regions.
Unfortunately, current attention accelerators still suffer from computation and memory access inefficiencies, as they fail to exploit bit-grained opportunities to eliminate the high-overhead predictor. Table I summarizes their features. The majority of existing works [ham20203, ham2021elsa, lu2021sanger, qu2022dota, yang2022dtatrans] focus on accelerating attention by alleviating computation overhead. For example, ELSA [ham2021elsa], Sanger [lu2021sanger], DOTA [qu2022dota], FACT [qin2023fact] adopt techniques like binary hashing, half-precision MSB, low-rank approximation, log-domain shifting to accelerate computation. However, these methods overlook memory optimization. While SpAtten [wang2021spatten]. SOFA [wang2024sofa] realizes this challenge, their coarse-grained strategies, like hybrid quantization, and cross-stage tiling, fail to exploit fine-grained, bit-level optimizations. Further, all current works rely on extra sparsity predictors, incurring substantial overhead. Notably, DTATrans [yang2022dtatrans] and SpAtten [wang2021spatten] guide sparsity using attention scores from the previous layer. While this strategy partially reduces predictor overhead, it necessitates resource-intensive retrain to recover accuracy. These limitations motivate us to design an efficient attention accelerator that jointly optimizes computation and memory at fine granularity, while eliminating the sparsity predictor.
IV Algorithm Optimizations of PADE
To effectively support the BSF strategy, we propose three key optimizations: BUI-GF, BS-OOE, and ISTA. BUI-GF ensures precise pruning in bit-wise operations, BS-OOE optimizes hardware utilization, and ISTA maintains sparsity while enhancing I/O efficiency via tiling attention.
Figure 6: Illustration of BUI with dot-product $\mathbf{Q}_{i}$ $\times$ $\mathbf{K}_{j}$.
IV-A BUI-enabled Guarded Filtering (BUI-GF)
We begin by analyzing the mathematical properties of the softmax function, revealing its intrinsic potential for a simple yet efficient max-based pruning decision. Without loss of generality, consider a two-element input vector $x_{0},x_{1}]$, where $x_{1}$ is the max element and $x_{1}$$x_{0}$$\Delta$. As depicted in Eq.(1), the softmax output for $x_{0}$ decays exponentially with the gap $\Delta$ from $x_{1}$. This indicates that an element’s contribution diminishes rapidly as its distance $\Delta$ from the max increases.
$x_{1}\!\!=\!\!x_{0}\!+\!\Delta\!\Rightarrow\!\mathrm{softmax}(x_{0})\!=\!\frac{e^{x_{0}}}{e^{x_{0}}\!+\!e^{x_{0}+\Delta}}\!=\!\frac{1}{1\!+\!e^{\Delta}}<\frac{1}{e^{\Delta}}\!\!$ (1)
Naturally, a straightforward way to combine BSF with a max-based pruning strategy is to estimate the attention via partial bit planes of Keys, identify the max value, then use it to define a pruning threshold. However, such a crude method will incur severe estimation error, as analyzed in §III-B.
To this end, we propose a bit-level uncertainty interval (BUI) enabled guarded filtering (BUI-GF). We first introduce the BUI, which quantifies the potential variation in a dot-product, i.e., $\mathbf{Q}_{i}\mathbf{K}_{j}$ caused by the remaining bit planes of $\mathbf{K}_{j}$. Specifically, for a $p$-bit integer $b_{p-1}b_{p-2}...b_{0}$ with 2’s complement format, its value $x$ is:
$x=-b_{p-1}2^{p-1}+\sum\nolimits_{i=0}^{p-2}b_{i}2^{i}.$ (2)
In this format, all bits except the sign bit ($b_{p-1}$) contribute a non-negative value, meaning that each additional bit can only increase or maintain the magnitude. Based on this property, Figs. 6 (a)(b) exemplify the BUI. In the example, $\mathbf{Q}_{i}$ denotes the $i$-th row of $\mathbf{Q}$ matrix, containing four entries with full 8-bit precision, while $\mathbf{K}_{j}$ is bit-serially processed: 1 bit in Fig. 6 (a) and 2 bits in Fig. 6 (b). For positive elements of $\mathbf{Q}_{i}$, BUI sets the unknown bits of $\mathbf{K}_{j}$ to 1 (shown in blue) and for negative entries, it set them to 0 (shown in red), yielding the potential largest score $S_{i,j}^{r,\max}$, as they account only for positive contributions. Here, $r\in$$[0,7]$ denotes the number of processed bit planes in 8-bit quantization. Conversely, BUI flips the unknown bits to obtain the potentially smallest score $S_{i,j}^{r,\min}$. We model this process as follows:
$S_{i,j}^{r,\min}=S_{i,j}^{r}+I_{i}^{r,\min},\hskip 11.38109ptS_{i,j}^{r,\max}=S_{i,j}^{r}+I_{i}^{r,\max},$ (3)
where $S_{i,j}^{r}$ denotes the conservative value by setting all unknown bits of $\mathbf{K}_{j}$ to zero, and $I_{i}^{r,\min}$ and $I_{i}^{r,\min}$ are the uncertainty intervals decided only by $\mathbf{Q}_{i}$. For example, in Fig. 6 (a), we only know the MSB of $\mathbf{K}_{j}$, and by applying Eq. (3), the BUI is determined to be the low bound (LB) $S_{i,j}^{0,\min}=-101.75$ and the upper bound (UB) $S_{i,j}^{0,\max}=84.25$.
Based on the BUI, we propose the BUI-GF strategy, which consists of two main steps, as illustrated in Fig. 7. First, the BUI-GF determines the pruning threshold using the Eq. (4):
$\mathcal{T}=\max(S_{i,:}^{:,\min})-\alpha\times radius,\hskip 5.69054pt0\leq\alpha\leq 1,$ (4)
Figure 7: Illustration of the BUI-GF strategy.
where $S_{i,:}^{:,\min}$ denotes the LBs of the estimated attention score for the $i$-th row. Notably, the scores are not confined to a specific bit plane, but are derived from all current processed bit planes. Based on our experiments, we set the default radius to 5 and use a parameter $\alpha\in$$[0,1]$ to control the threshold $\mathcal{T}$. By adjusting the $\alpha$, we can control the pruning ratio (DSE see §VI-D). Then, the BUI-GF compares the UBs of attention scores with the threshold, retaining the Ks (e.g. 0,3,4,6,7), whose attention values are greater than this threshold.
IV-B BS-OOE: Improving Hardware Utilization
The BUI-GF strategy enables accurate pruning decisions based on partial scores, which are incrementally estimated using bit-wise planes of the Key tensor. However, as analyzed in §III-B, bit-grained execution will lead to hardware underutilization. To this end, BS-OOE employs a two-pronged approach: First, a bidirectional sparsity (BS) ensures load balancing across PEs. Second, bit-wise out-of-order execution (OOE) hides DRAM access latency.
The core idea of BS is that 1 is also a form of sparsity. Without loss of generality, the dot-product between an $N$-element Query and Key can be formulated as:
$\sum\nolimits_{j=0}^{N-1}q_{j}k_{j}=\sum\nolimits_{b=0}^{p-1}2^{b}\times\sum\nolimits_{j=0}^{N-1}q_{j}\times k_{j}^{b},$ (5)
where $k_{j}^{b}$ is the $b$-th bit of element $k_{j}$. Since each bit of $k_{j}$ can only be either 0 or 1, the second partial sum on the right-hand side of Eq. (5) can be reorganized as:
$\sum\nolimits_{j=0}^{N-1}\!\!q_{j}\times k_{j}^{b}=\!\!\!\sum\nolimits_{\forall j:k_{j}^{b}=1}\!q_{j}=\!\!\sum\nolimits_{j=0}^{N-1}\!\!\!q_{j}\!-\!\!\sum\nolimits_{\forall j:k_{j}^{b}=0}q_{j}.$ (6)
From Eqs. (5) (6), we observe that instead of accumulating the Query entries corresponding to bit-1, one can equivalently subtract the Query entries corresponding to bit-0 from the total sum of all Query entries. This transformation ensures that at most $50\%$ of the bits are involved in computation, thus effectively improving the load balance across PEs.
[…]
Figure 8: (a)(b) Bit-wise OOE execution, (c)-(e) benefits of BS-OOE strategy.
Figs. 8 (c)-(e) highlight the advantages of BS-OOE over naive bit sparsity + bit serial processing, by comparing the timeline of PE0 in the scenarios depicted in Figs. 8 (a)(b). As shown in Fig. 8 (c), naive bit sparsity processing suffers from workload imbalance, causing severe computation time discrepancies across PEs. This results in scattered DRAM accesses, idle cycles, and high latency. In Fig. 8 (d), BS alleviates this imbalance between PEs, enabling memory request merging and reducing row activation during DRAM access. However, computation resources remain underutilized during memory access. In contrast, as depicted in Fig. 8 (e), BS-OOE further leverages bit-wise out-of-order execution to improve resource utilization. Specifically, while PE0 waits for DRAM to return $K_{0}^{1}$, it continues to process $K_{1}^{0},..\,,K_{4}^{0}$, without idling.
IV-C ISTA: Enhancing IO Efficiency
Reducing unnecessary data movement is critical for supporting ultra-long sequences. While tiling (e.g., FlashAttention) offers a promising solution, it is incompatible with sparse attention because it breaks the row-wise dependency of softmax, on which BUI-GF (§IV-A) relies for max-based pruning.
Figure 9: Tile-level sparsity decision.
To address this, we re-examine softmax properties and show that pruning decisions can be performed on subsets with proper adjustments. Eq. (7) shows that the softmax denominator grows monotonically as more elements are added, due to the non-negativity of exponentials. Therefore, if a token’s score falls below the threshold within a subset, its global row-wise softmax score can only be lower. This enables the safe application of the BUI-GF strategy within tiled regions.
$softmax(x_{i})=\frac{\exp(x_{i})}{\sum_{j=0}^{N-1}\exp(x_{j})}\leq\frac{\exp(x_{i})}{\sum_{j\in subset}\exp(x_{j})}$ (7)
To identify retained keys within a subset, a key is retained if all its bit planes are processed and it remains unpruned. As depicted in Fig. 9, at cycle 0, only the partial score of $K_{0}$ is calculated, with the subset observation window size initialized to 1. By cycle 15, scores relative to six Keys have been computed, expanding the window to size 6. During each bit plane calculation, the BUI-GF works continuously to check if pruning. If a Key (i.e., $K_{0}$) remains unpruned after processing its least significant bit (LSB) plane, it is deemed an important token, and its index is stored in the Retained Key Board.
Driven by the tile-level sparsity decisions, the ISTA algorithm is detailed in Fig. 10 (c). Its process begins by performing sparse QK computations, retaining and storing the indices (IDs) and scores (SIDs) of selected Keys (lines 4-6). Once the number of retained entries reaches the tile size $B_{c}$ (line 3), the corresponding Vs are fetched for subsequent computation (line 8). As more tiles are processed, the maximum value (line 9) and the exponential sum (line 11) are progressively updated and propagated across tiles. Finally, the attention output is computed based on the accumulated results (line 13).
However, a key observation is that naive left-to-right computation causes frequent updates of the maximum across tiles (line 9 in Fig. 10 (c)), leading to a series of redundant operations (line 11). Specifically, each time the maximum is updated, it triggers one subtraction, one exponentiation and two multiplications of scalar and vector (see highlights in lines 11-12). We employ the arithmetic complexity model [brent2010modern, wang2021efficientMIMO, wang2023efficient] to normalize the complexity for different operations. As profiled in Fig. 10 (b), for SS=$2048$ and $B_{c}$=16, this leads to about 30% increase in computation complexity compared to the vanilla implementation. Moreover, the overhead becomes more as $B_{c}$ decreases, due to more frequent updates.
Figure 10: (a) Illustration for head-tail interleaving updating. (b) Heavy updating cost across tiles. (c) Detailed procedure of the ISTA algorithm.
Figure 11: (a) Overview of the PADE architecture. (b) Bit-wise PE lane. (c) BUI Generator. (d) BUI-GF Module. (e) Decision Unit.
To minimize the overhead associated with the ‘max updating’, it is advantageous to prioritize dominant tokens. Note that the reduction in operations refers to the redundant computations shown in lines 11–12, rather than the comparison operations in line 9. To this end, we propose a head-tail interleaved updating strategy that exploits the locality property of attention: Recently generated tokens and the initial token typically exhibit higher weights than others [tian2023scan, jiang2024minference]. As illustrated in Fig. 10(a), the update begins with the initial region, then jumps to the recent region, and subsequently returns to the post-initial region, repeating this interleaved pattern across tiles. This pattern reduces unnecessary maximum value updates, resulting in a $20\%$-$40\%$ reduction in total operations. It is important to note that, without attention locality, the performance of head-tail interleaving is on par with regular execution and not worse.
V Architecture and Hardware Innovation
V-A Architecture Overview
Fig. 11 (a) depicts the overall architecture of the PADE, which incorporates two major components:
1) Query-Key Processing Unit (QK-PU): This unit computes the dot-product of the Query and Key matrices (i.e., $\mathbf{QK}^{T}$). It includes a PE array, eight BUI-GF modules, a Q_\_sum generator, a sparsity scheduler, and a BUI generator. The PE array comprises eight rows, each containing 16 bit-wise PE lanes, dedicated to processing a single query. Together, these components support both the BUI-GF and BS-OOE strategies.
[…]
V-B Overall Dataflow
In PADE, self-attention operands use 8-bit precision, with each Key vector divided into eight 1-bit planes. During the prefill stage, PADE processes eight queries within a head, whereas during the decoding stage, it handles different queries across multiple heads. Its detailed process is as follows (Fig. 11 (a)):
First (❶), before the $\mathbf{Q}_{i}\mathbf{K}^{T}$ computation, the BUI Generator initializes eight uncertainty interval pairs $(I_{i}^{r,\min},I_{i}^{r,\max})$, where $r\in$$[0,7]$, based on the input $\mathbf{Q}_{i}$. Each pair corresponds to a specific bit plane. These uncertainty interval pairs are then stored in a lookup table (LUT), as shown in Fig. 11 (c). Next (❷), $16$ PE lanes in a PE row perform the dot product for $\mathbf{Q}_{i}\mathbf{K}^{T}$ and in an out-of-order manner in parallel. Following this (❸), the BUI-GF Module, as depicted in Fig. 11 (d), calculates the pruning threshold $\mathcal{T}_{i}$ using the max value among the all current score $S_{i}^{:,\min}$ with the BUI-GF logic (Eq. (4)). Finally (❹), the pruning threshold $\mathcal{T}_{i}$ is broadcast to all PE lanes in a row, enabling the evaluation of whether the score of the token $j$ satisfying $S_{i,j}^{r,\max}$$\mathcal{T}_{i}$, as depicted in Fig. 11 (e). If true, the PE lane requests the next bit plane for further computation. Otherwise, the token $j$ is immediately pruned. This process is repeated until the LSB is reached. The remaining scores are sent to V-PU to produce final outputs.
V-C ScoreBoard-Based Result Reusable PE Lane
To perform bit-serial speculation for $\mathbf{QK}$, a straightforward solution is to incrementally compute over the bit planes of $\mathbf{K}$. Specifically, it accesses one bit plane (MSB) in the first round, two bit planes (MSB, MSB-1) in the second round, and so on, until the final round. However, repeated memory accesses in this scheme lead to significant power consumption.
To enable result reuse, we dedicate a scoreboard-assisted bit-wise PE lane, as shown in Fig. 11 (b), which comprises three key components: a grouped lightweight sparsity ANDer tree (GSAT) for 64-input dot-product and two modules supporting BUI-GF for out-of-order execution. 1) To avoid repeatedly loading bit planes, each PE lane integrates a scoreboard that temporarily caches partial scores $S_{i,j}^{r}$ for unpruned tokens. 2) The Decision Unit performs BUI-GF logic, makes pruning decisions and selects the next bit plane to fetch.
These modules operate collaboratively to enable efficient early pruning while minimizing redundant memory access. Initially, the GSAT computes the partial dot product $\Delta S_{i,j}^{r}$ from an 8-bit vector $\mathbf{Q}_{i}$ and a 1-bit Key plane $\mathbf{K}_{j}^{r}$. Meanwhile, the scoreboard is accessed via the Key’s index $j$. If a previous partial score $S_{i,j}^{r-1}$ exists, it is fetched and updated as $S_{i,j}^{r}$$S_{i,j}^{r-1}$ ++ $\Delta S_{i,j}^{r}$. Otherwise, indicated that the current bit plane is MSB, $\Delta S_{i,j}^{r}$ is directly written to the scoreboard, and Hit signal is pulled down to show no prior score is available.
Decision Unit. The unit determines pruning by receiving the max uncertainty interval $I_{i}^{r,\max}$ and the partial score $S_{i,j}^{r}$. It checks whether $S_{i,j}^{r}$$I_{i}^{r,\max}$$\mathcal{T}_{i}$ holds. If true, it requests the next bit plane of $\mathbf{K}_{j}$, i.e., $\mathbf{K}_{j}^{r+1}$, and updates the partial score in the scoreboard. Otherwise, it evicts the token entry from the scoreboard and requests the next Key vector from DRAM.
BUI-GF Module. As shown in Fig. 11 (d), the BUI-GF module reads scores from registers and adds the min uncertainty interval $I_{i}^{r,\min}$ to compute the lower bound of scores. Based on these values and a predefined ratio $\alpha$, the Threshold Updating Module applies BUI-GF logic (Eq. (4)) to generate threshold $\mathcal{T}_{i}$, which is then broadcast to all PE lanes in rowi.
V-D Grouped Lightweight Sparsity ANDer Tree (GSAT)
To exploit bit sparsity, a direct design is to select and accumulate activations corresponding to non-zero bits. However, this requires large multiplexers, reducing efficiency. Specifically, to perform a 64-dimensional dot product, such a coarse design will require at least 32 64-input multiplexers.
Given BS guarantees at least $50\%$ sparsity in a bit-vector of arbitrary length, it is possible to reduce the MUX cost with a smaller group size. Based on this insight, we propose a grouped lightweight ANDer tree architecture. As depicted in Fig. 11 (b), for a 64-input dot-product, we first decompose it into eight accumulation sub-groups, each comprising 8 dimensions. For each sub-group, in the worst-case scenario, the selected query elements within the sub-group $\{q_{0},\cdots,q_{7}\}$ will be $\{q_{4},\cdots,q_{7}\}$. Therefore, in a sub group, only four 5:1 multiplexers are required: the first multiplexer selects among $\{q_{0},\cdots,q_{4}\}$, the second among $\{q_{1},\cdots,q_{5}\}$, and so on. Compared to the naive design, MUX cost is reduced (32×64:1 $\to$ 32×5:1), with a trade-off of more subtractors.
Insight: There is an optimal sub-group size that minimizes hardware overhead. A smaller sub-group size reduces multiplexer overhead but adds extra cost from subtractors and Q_sum generators. The DSE is provided in Fig. 17 (a).
BS Scheduler. PADE adopts a low-cost BS scheduler to manage the operations within each PE, as depicted in Fig. 12. To control the bit-serial dot product, the scheduler first identifies whether a bit plane of the Key vector contains more zeros or ones. It then sends the original or flipped bit column to a priority encoder. The priority encoder operates on five consecutive bits of the plane at a time. For example, at the first time, it receives $k_{0}$$k_{4}$}, followed by $k_{1}$$k_{5}$} in the next, and so on. The encoder detects the location of the first “1” bit within each 5-bit vector. If such a bit exists, it is masked, and the remaining bits are propagated to the next time step. Otherwise, if the vector contains only zeros, the encoder asserts val=0 to disable the corresponding bit-serial multiplier in the PE, as depicted in Fig. 11 (b).
Figure 12: Detailed architecture of bidirectional sparsity (BS) scheduler.
V-E Reuse-Aware Reorder Scheduling (RARS)
Due to attention sparsity, the remaining scores are distributed irregularly. For example, in Fig. 13 (a), the 0-th score row S0 retains elements at positions 0-3, which are multiplied with the corresponding V0-V3 vectors (refer to Fig. 3).
However, a naive design leads to redundant V-vector memory accesses during the computation of $\mathbf{S}\times\mathbf{V}$. As shown in Fig. 13 (a), assuming each PE row in V-PU processes two V vectors per round, a naive left-to-right execution computes the dot products (S0,V0V1), (S1,V2V3), (S2,V4V5), (S3,V2V3) in round 0, followed by (S0,V2V3), (S1,V4V7), (S2,V6V7), (S3,V4V7) during round 1. Due to data reuse inefficiency, some shared V vectors must be reloaded, resulting in a total of 11 V-vector memory accesses.
To implement the ISTA strategy efficiently, we propose reuse-aware reorder scheduling (RARS) to reduce redundant memory access. As shown Fig. 13 (d), V2 and V3 are shared among three scores: S0, S1 and S3, making them are prioritized for initial scheduling. Then, RARS selects V5 and V6, which are used exclusively by the remaining score S2. As a result, V2, V3, V5 and V6 are grouped for execution in round 0. Such greedy search continues until all scores are allocated adequate Vs. As depicted in Fig. 13 (e), compared to the default left-to-right computation order, RARS reduces $30\%$ memory access.
We design an efficient scheduler to implement RARS. As shown in Fig. 13 (c), condition statements and control logic are handled by an FSM controller. A single-port read-write ID buffer, indexed by score-derived bitmasks, stores the corresponding V-vector IDs. For example, V5 and V6, used exclusively by S2, are stored in buffer-0010. Guided by RARS logic, the FSM retrieves the buffer entries and dispatches them to the issuing FIFO in an optimized execution order.
VI Evaluation
VI-A Experimental Setup
Baseline comparisons: We compare PADE with five SOTA attention accelerators: Sanger [lu2021sanger], Spatten [wang2021spatten], Energon [zhou2022energon], DOTA [qu2022dota], SOFA [wang2024sofa]. For fair comparison, all designs are normalized to a 28nm process and evaluated under identical conditions: PE arrays occupy the same area as PADE and work in 800 MHz, on-chip SRAM is set to 352kB, and peak HBM bandwidth is 256 GB/s, with 4 pj/bit [o2017fine].
Figure 13: Reuse-aware reorder scheduler (RARS).
TABLE II: Accuracy of Different Transformer Models with MXINT8, FP16, INT8 and PADE Configurations (S: Standard, A: Aggressive).00footnotetext:
Model LlaMa2-7B LlaMa3-8B OPT1B3 Bloom1B7 Qwen7B ViT-L/16 PVT
Task Dolly Wikili. MBPP Wiki2 MMLU Winog. Dolly Wikili. MBPP Wiki2 MMLU Winog. Wikili. MBPP Wikili. MBPP Wikili. MBPP Image VTAB Image VTAB
MXINT8 36.5 39.3 17.5% 5.63 35.2% 69.8% 40.9 43.6 23.3% 5.01 42.2% 75.1% 36.1 11.9% 44.6 16.3% 46.8 30.5% 85.5% 72.8% 89.7% 77.5%
FP16 36.4 39.1 17.5% 5.71 35.1% 69.4% 40.8 42.7 21.8% 5.11 41.2% 74.2% 36.2 11.9% 44.3 16% 46.6 30% 85.3% 72.7% 89.4% 77.3%
INT8 36.4 38.9 17.2% 5.73 34.7% 69.3% 40.7 42.7 21.6% 5.13 40.9% 73.7% 35.9 11.6% 44.1 15.7% 46.4 29.2% 85.3% 72.5% 89.3% 77.1%
PADE (S) 36.3 38.9 17.2% 5.75 34.6% 69.2% 40.6 42.6 21.5% 5.13 40.7% 73.7% 35.9 11.5% 44.0 15.6% 46.3 29.2% 85.3% 72.5% 89.3% 77.1%
PADE (A) 36.1 38.4 16.5% 5.80 34.1% 68.7% 40.5 42.0 21.0% 5.19 40.2% 72.8% 35.3 11.0% 43.6 15.2% 45.9 28.4% 84.9% 72.4% 89.1% 76.8%
MMLU, WinoGrande, MBPP, Imagenet, VTAB are evaluated by accuracy. Dolly, Wikilingua are evaluated by ROUGE-1. Wikitext2 is evaluated by PPL, where lower is better.
Figure 14: Normalized computation and memory access across diverse Transformer models and tasks. Spatten* performs additional fine-tuning.
Benchmarks: We evaluate PADE on several representative Transformer models across NLP and CV tasks. For NLP tasks, Llama2-7B [touvron2023llama2], Llama3-8B [grattafiori2024llama], Qwen7B [bai2023qwen], Bloom1B7 [le2022bloom] and OPT1B3 [zhang2022opt], for six tasks.These tasks include language modeling (Wikitext-2 (S=2k) [merity2016pointer], Wikilingua (S=2k) [ladhak-wiki-2020], Winogrande (S=0.25k) [sakaguchi2021winogrande]), language understanding (MMLU, S=0.5k) [hendrycks2020measuring], code generation MBPP (S=1k) [austin2021program], long context processing dolly (S=15k) [conover2023free]. For CV tasks, we choose ViT-L/16 (S=576) [dosovitskiy2020image] and PVT (S=3k) [wang2021pyramid] on ImageNet-1k [deng2009imagenet] and JFT [sun2017revisiting] classification.
Quantization Accuracy. All pre-trained models are sourced from Pytorch [paszke2017automatic] and HuggingFace [wolf2020transformers]. INT8 baselines derived via post-training quantization, where only the weights and activations (QKVs) are quantized to INT8, while non-linear operators (e.g., softmax) remain in FP16 precision. As shown in Table. II, the INT8 baseline incurs less than a $1\%$ average accuracy drop from FP16, thus confirming its validity.
Hardware Evaluation: Table III lists the hardware configuration of PADE. We implement the RTL design for PADE and utilize Synopsys DC on TSMC 28nm CMOS technology to estimate the logic area and power. The power, area, and read/write bandwidth of on-chip SRAM buffers are estimated through CACTI [muralimanohar2009cacti]. Off-chip HBM modeling involves simulating access patterns and row activation under various data layouts (Fig. 22), capturing HBM’s burst behavior. We derive memory latency from Ramulator [kim2015ramulator], and estimate IO power following the methodology in [cavigelli2016origami, andri2016yodann, wang2017energy]. We extract each stage’s cycles by simulating the RTL with Verilator [snyder2004verilator], and use a custom cycle-level simulator to evaluate performance.
TABLE III: On-chip Hardware and Off-chip DRAM Configurations of PADE
On-chip Buffer 320KB SRAM for Key and Value buffers;
32KB Q buffer
QK-PU 128 Bit-wise PE Lanes; A Q_\_sum Generator;
8 BUI-GF modules; A BUI Generator; A Sparsity Scheduler
- Bit-wise PE Lane 64-dim $\times$ 8-bit $\times$ 1-bit Grouped Sparsity ANDer tree;
32 entry $\times$ 45-bit Scoreboard
V-PU A Systolic Array with 8 $\times$ 16 INT8 PEs;
A 128-input FP16 APM; A RARS Scheduler
[…]
GPU comparison We benchmark on an Nvidia H100 using SOTA TensorRT-LLM [Nvidia2023] with FlashAttention3 [shah2024flashattention]. To exclude the software overhead, we measure execution time with cudaEvent, isolating GPU execution from CPU interference. The GPU is dedicated during testing, and large batch sizes are used to amortize data transfer costs. Non-computational phases are excluded using nvprof. Power is measured via nvidia-smi; dynamic power is computed as the difference between active and idle states. Each experiment is run 2k times, discarding the top and bottom $15\%$ before averaging.
GPU test configurations: We test all different datasets with their allowed sequence lengths ranging from 0.25k to 15k. For example, MMLU (0.5k) and Dolly (15k). We measure the total inference latency, including the prefill and decoding. Specific prefill and decoding lengths are determined by the dataset itself. For batch size, we will select the configuration from [8, 128] that maximizes GPU computational efficiency based on the sequence length of each dataset.
Figure 15: (a)(b) Accuracy comparison with current sparse attention methods. (c) Speedup and efficiency gain comparison.
VI-B Algorithm Performance
Algorithm settings: INT8 models serve as the accuracy baseline, and $\alpha$ (Eq.(4)) is adjusted in 0.1 increments to evaluate the accuracy and overhead for each benchmark. Two PADE configurations are evaluated: standard ($0\%$ loss), aggressive ($1\%$ loss).
Fig. 15 (a)(b) compares the accuracy of three representative software-only sparse attention methods, two predictor-free works (SpAtten [wang2021spatten], DTATrans [yang2022dtatrans]), and PADE. The Sparsity Level denotes the ratio between the computation cost of sparse execution (prediction + actual computation) and that of dense execution. StreamingLLM [xiao2023efficient] adopts a static sparsity pattern—retaining only the initial and recent tokens, while MInference [jiang2024minference], and DoubleSparsity [yang2024post] rely on runtime sparsity prediction. Four key observations are made: 1) StreamingLLM performs the worst, due to its lack of adaptivity in capturing input-dependent sparsity patterns. 2) MInference improves accuracy by combining dynamic prediction with predefined sparsity patterns but remains limited by restricted pattern diversity. 3) DoubleSparsity introduces a more flexible dynamic sparsity mechanism and integrates channel sparsity to reduce prediction overhead. However, as its prediction computation and memory access cannot be reused in subsequent steps, it suffers from inefficiency, yielding slightly lower accuracy than PADE at the same sparsity level. 4) For SpAtten and DTATrans, comparable accuracy to PADE can be achieved only after fine-tuning. This is mainly because both methods eliminate the predictor by using the previous layer’s attention distribution to guide pruning in the current layer, which introduces significant errors without fine-tuning. In contrast, PADE consistently achieves the best performance across all settings, benefiting from its fine-grained sparsity prediction and efficient reuse of both computation and memory access.
[…]
Figure 16: (a) Latency reduction for BUI-GF, BS-OOE and ISTA. (b) Exploring the trade-off between accuracy and sparsity.
Computation Reduction. Fig. 14 compares the computation reduction across accelerators with $0\%$ accuracy loss. Spatten, without retraining, yields the lowest reduction and serves as the baseline. Energon enhances performance via progressive precision prediction, achieving a $32\%$ reduction, and outperforming coarse-grained prediction accelerators like Sanger and DOTA. However, it lacks computation reuse. SOFA mitigates this by using log-domain differential leading-one computation, but still relies on an added predictor, limiting its reduction to $45\%$. In contrast, PADE eliminates the need for a predictor by leveraging fine-grained bit-level early termination and bit reuse, achieving a reduction of up to $71.6\%$.
Memory Access Reduction. As shown in Fig. 14, Sanger, which employs an extra 4-bit MSB predictor, serves as the baseline. DOTA adopts low-rank approximation but fails to mitigate prediction bitwidth overhead. Energon and Spatten* (with finetune) partially alleviate this via mixed precision, but still rely on an extra predictor, limiting their reduction to $21\%$ and $42\%$, respectively. SOFA reduces memory overhead via cross-stage tiling yet remains constrained by the extra predictor. In contrast, PADE achieves an average memory reduction of $75.8\%$ across both long- and short-sequence tasks, attributed to the BSF strategy. When compared to Spatten without fine-tune, PADE achieves $3.4\times$ higher memory reduction.
VI-C Design of Architecture Parameters
The key architectural parameters in PADE are decided by detailed workload profiling and DSE exploration.
Figure 17: Exploration of optimal (a) Sub-group size. (b) Scoreboard entry.
(1) Computational throughput: The INT-8 throughput ratio between QK-PU and V-PU is set to 8:1 (see Table III), derived from the typical QK-to-SV computation ratio observed in LLM workloads due to sparsity.
(2) Buffer Sizes: A 320 KB KV buffer and a 32 KB query buffer are allocated, sufficient for 12.8k tokens under typical sparsity ratio of 0.2 (64-d embedding) and 256 queries, respectively. These configurations balance on-chip storage efficiency and multi-phase processing for longer sequences.
(3) Optimal Sub-group Size. We conduct a DSE to determine the optimal PE group size (Recall §V-D). As shown in Fig. 17 (a), a sub-group size of 8 minimizes area and power overhead, which is therefore adopted in our PADE accelerator.
(4) Optimal Scoreboard Size. A larger scoreboard can cache more partial sums, which improves PE utilization but increases area overhead. Fig. 17 (b) shows that PE utilization saturates at around 32 entries. This indicates that, at this point, memory access and computation rates are balanced, preventing PE computation from being blocked. Therefore, PADE adopts a 32-entry scoreboard, as in Table III.
VI-D Architecture Evaluation
Ablation. We conduct an ablation study to evaluate the latency reduction of BUI-GF, BS-OOE, and ISTA against a baseline dense attention accelerator derived from PADE, but without sparse processing modules. As shown in Fig. 16 (a), BUI-GF reduces average latency by $30\%$, mainly by predictor-free token sparsity. Further, BS-OOE realizes $24\%$ latency reduction via improved hardware utilization. Finally, ISTA achieves $27\%$ decrease, via efficient tiling and data reuse.
Figure 18: (a) Breakdown for the overhead of bit shifting. (b) Latency and Efficiency gain of PADE over Nvidia H100 GPU.
Accuracy &\& Sparsity Trade-off. The BUI-GF may affect accuracy, as it introduces a parameter $\alpha$ to pruning KVs (§IV-A). Fig. 16 (b) shows the impact of $\alpha$ on accuracy and sparsity using LLaMA2-7B on MMLU (reasoning) and MBPP (generation). Overall, a smaller $\alpha$ results in more aggressive pruning, decreasing accuracy but increasing sparsity. There are some key observations: For generation tasks (MBPP), accuracy drops noticeably when $\alpha$<<$0.6$. In contrast, for reasoning tasks (MMLU), the model is more tolerant to pruning, with accuracy degrading evidently only when $\alpha$<<$0.5$. This may be because reasoning tasks rely on vital tokens for inference, resulting in higher token redundancy. On the other hand, the sparsity gains begin to diminish when $\alpha$<<$0.5$, likely due to overly aggressive pruning, which harms crucial tokens and limits further sparsity. Therefore, to strike a balanced trade-off between accuracy and sparsity, we empirically set $\alpha$ within the range of 0.5–0.6.
[…]
Figure 19: Energy efficiency and throughput gain breakdown.
Efficiency gain breakdown. Fig. 19 (a) gives the energy efficiency breakdown. With a dedicated ASIC and customized datapath, PADE achieves a $4.0\times$ gain over the GPU baseline. While the token sparsity-leveraged BUI-GF theoretically reduces computation by $3.7\times$, practical efficiency improves by only $1.4\times$. This is due to redundantly loading repeated bit planes. After adding the dedicated scoreboard-based result reusable PE lane, the performance jumps by $2.2\times$. Similarly, directly applying the BS-OOE scheme and tiling-target ISTA scheme yields only $1.58\times$ and $1.43\times$ efficiency gain. This is due to mismatched computational granularity and the presence of unused V vectors, which ultimately lead to severe resource underutilization. By contrast, deploying tailored engines can further bring $2.07\times$ and $1.69\times$ efficiency gain effects.
Area and Power. Fig. 20 presents the area and power breakdown of PADE. Occupying 4.53 mm2 and consuming 591 mW, PADE achieves a peak energy efficiency of 11.36 TOPS/W. The added BUI Generator and BUI-GF modules adaptively respond to attention distribution for token pruning, incurring only $4.9\%$ area and $12.1\%$ power overhead. Also, integrating the Scoreboard and Decision Unit into PE lanes enables stage fusion with just $5.8\%$ area and $4.9\%$ power cost. Despite the modest overhead, eliminating the sparsity predictor and reducing off-chip memory access yield notable speedup and efficiency gains. As depicted in Fig.19 (b), compared to baseline without sparse processing modules, software-hardware co-design BUI-GF, BS-OOE and ISTA bring $2.12\times$, $1.9\times$ and $1.3\times$ throughput gain, respectively. In summary, PADE represents a deliberate tradeoff, achieving substantial efficiency gains with minimal resource overhead.
Data Layout. In PADE, the data layout is carefully co-designed across the off-chip DRAM and on-chip SRAM to maximize memory bandwidth utilization. As depicted in Fig. 22 above, K is bank-interleaved along the bit dimension, meaning that each DRAM bank stores a distinct bit-plane of the K tensor, enabling efficient bit-plane access when performing bit-level computations. In contrast, Q and V are bank-interleaved along the hidden (H) dimension such that 8-bit data is read continuously. When the data are fetched into the on-chip SRAM, the layout is reorganized to match the PE access pattern. In the Q,V SRAM, each row primarily stores different bits of the same element, whereas in the K SRAM, each row stores the same bit plane (e.g., MSB) from multiple elements across the hidden dimension.
Figure 20: Area/Power of PADE at TSMC 28nm, 800MHz.
VI-E Comparison to SOTA Accelerators
[…]
Figure 21: Speedup and energy breakdown comparisons.
Figure 22: DRAM, SRAM layouts and overall dataflow of PADE.
Workload Balance. Fig. 23 (a) presents a detailed breakdown of execution time across varying numbers of PE lanes to illustrate load balance effects. To better show PADE’s advantages, we compare it against a SOTA bit accelerator, BitWave [shi2024bitwave], which leverages bit-flipping to enhance bit-plane sparsity. Since each PE lane integrates multiple bit-serial multipliers, intra-PE stall arises when certain multipliers handle more effective bits. Inter-PE install, in contrast, results from variations in bit sparsity across different key vectors. As the number of PE lanes increase, BitWave suffers from greater intra- and inter-PE imbalance, as it only exploits bit-0 sparsity, which exhibits large variability. In comparison, PADE adopts a more balanced bit sparsity distribution, thereby achieving around $30\%$ higher PE utilization.
Figure 23: (a) Breakdown of execution cycles versus the number of PE lanes. (b) Bandwidth utilization, DRAM access, and speedup comparisons.
[…]
VI-F Discussion for Deployment and Scalability
System Integration. PADE functions as a co-processor working collaboratively with the GPU. The GPU handles dense computations such as QKV projection and FFN, while PADE accelerates attention via sparsity. As depicted in Fig. 24 (a), both processors execute instructions issued by the host CPU and share the device memory, enabling direct data exchange without additional transfers. During the K generation, GPU executes an extra data conversion operation to store K in HBM with a bit-plane-first layout, as shown in Fig. 22.
[…]
Figure 24: Conceptual integration of PADE with a GPU-like NN accelerator.
Extension for MXINT. The micro-scale format performs fine-grained quantization along the channel dimension by grouping data into 32-element segments [rouhani2023microscaling]. PADE ensures compatibility by applying group-wise scaling to the bit uncertain interval (BUI) (§IV-A). As shown in Fig. 25 (a), when processing 64-length Q and K vectors, the micro-format partitions them into two 32-element groups and quantizes each group using calibration-derived factors ($\Delta$), resulting in group-wise BUI scaling. As depicted in Fig. 25 (b), PADE achieves compatibility via two steps. ❶ First, it performs bitwise serial multiplication within each 32-length group, with the BUI derived from the strategy in §IV-A. The resulting BUI is then expanded using the quantization factor, eg., $\Delta_{\rm Q1}\Delta_{\rm K1}/\Delta_{\rm A}$. ❷ Finally, the max and min BUI values across all groups are aggregated to compute the overall BUI. This method can be extended to dot products of arbitrary length.
Extension for FP formats. Prior studies [sheng2023flexgen, xiao2023smoothquant, zhao2024atom] have demonstrated that the K and V tensors are highly amenable to quantization because the softmax normalization in self-attention naturally suppresses quantization noise. This property enables safe quantization to INT8 or even INT4 with negligible accuracy degradation. Motivated by this, when queries operate in FP format, PADE converts the INT-FP computation into a bit-serial form through exponent alignment, following methodologies adopted in prior works [chen2025bitmod, jang2024figna, fang2025anda]. The resulting bit-serial execution is fully compatible with PADE’s existing processing mechanism.
Diverse Quantizations. To assess PADE’s adaptability across quantization strategies and bit-widths, we extend both PADE and SOFA to QAT/PTQ INT8 and INT4 scenarios, as shown in Fig.26 (a). There are two key observations: (1) Compared to PTQ8, QAT8 leads to an  $6\%$ increase in energy consumption for SOFA. This is due to the more uniform data distribution under QAT, which reduces sparsity and renders SOFA’s predictor largely ineffective. In contrast, PADE also shows a slight energy increase, but the overhead remains negligible, as it avoids the use of any explicit predictor. (2) When the bit-width is reduced to 4 bits, SOFA’s energy efficiency gain drops significantly, as the predictor becomes the dominant source of power consumption under low-precision settings. In contrast, PADE’s energy gain decreases by only $2\%$, thanks to its predictor-free design that avoids this overhead.
Figure 25: Compatibility of the BUI-GF mechanism with the MX format.
Ultra-long Sequence Decoding. In long-sequence decoding, memory access optimization is critical due to the lack of data reuse, making it a key indicator of a design’s memory efficiency. As shown in Fig. 26 (b), in decoding, DRAM access accounts for over $85\%$ of total power overhead across all designs, due to the autoregressive nature of the workload. Notably, compared to the dense version, SOFA’s energy consumption increases significantly with sequence length—rising by nearly $40\%$ from 4K to 16K tokens. This is because SOFA must load all key vectors corresponding to the predicted sequence at each decoding step, making the predictor’s overhead grow rapidly. In contrast, PADE shows only a modest increase of about $5\%$ over the same sequence range, thanks to its predictor-free architecture. These results collectively demonstrate the advantages of PADE across diverse scenarios.
VI-G Limitations and Future Direction
While PADE significantly advances sparse-attention acceleration by eliminating the sparsity predictor through unified bit-serial stage fusion, several limitations remain for future works: (1) As the sequence length of large language models (LLMs) continues to scale, the need for distributed attention becomes inevitable. Extending PADE to distributed scenarios, especially in emerging wafer-scale architectures [wang2026WATOS, wang2026TEMP, wang2024tmac, he2025waferllm, hu2024wafer, tang2025moentwine], remains an open problem. (2) Beyond single-bit granularity, exploring multi-bit stage fusion may provide a more favorable efficiency–accuracy trade-off, which represents a promising direction for future PADE enhancements.
Figure 26: Efficiency with (a) diverse quantizations, (b) long sequence decoding.
VII Related Works
Efficient Attention Accelerators. Numerous attention accelerators [wang2025bitstopper, wang2025lapa, wang2025beta, li2020ftrans, ham20203, ham2021elsa, qu2022dota, fan2023taskfusion, yang2022dtatrans, hong2022dfx, liu2024hsconn, wang2024sofa, wang2022energy, fang2022algorithm, yazdanbakhsh2022sparse, li2022accelerating, lu2021sanger, wang2021spatten, zhou2022energon, qin2023fact, shen2022salo, fan2022adaptable, bai2024swat, zhao2024hardware, tuli2023acceltran, you2023vitcod, shen2022salo] have been proposed. Early works, such as A3 [ham20203] and ELSA [ham2021elsa] accelerate computation via approximation techniques. Recent efforts have shifted to jointly optimizing computation and memory. Energon [zhou2022energon] and SOFA [wang2024sofa] adopt fine-grained filtering and FlashAttention-optimized tiling techniques to alleviate memory overhead. However, they still rely on extra sparsity predictors, which become de facto latency and power bottlenecks after sparsification. PADE is the first to explicitly identify and address this issue. It removes the need for external predictors by fine-grained, bit-serial computation, while leveraging efficient bit-level early termination and reuse.
Neural network (NN) accelerator with sparsity. Numerous accelerators [wang2021efficient, hojabr2021spaghetti, wang2018low, gondimalla2019sparten, asgari2020alrescha, deng2021gospa, qin2020sigma, gudaparthi2022candles, fang2022algorithm, hanson2022cascading, lew2022anticipating, li2022ristretto, li2021escalate, li2022accelerating, liu2022s2ta, pavon2021via, rucker2021capstan, sadi2019efficient, walia2021fast, yang2020procrustes, Gon2023Eureka, wu2022sparseloop, wu2023highlight, shin2022griffin] exploit sparsity to accelerate NN inference. General-purpose sparse tensor accelerators [srivastava2020matraptor, mahmoud2020tensordash, kwon2019tensordimm, hegde2019extensor, kanellopoulos2019smash, chen2020tpspmv] support operations on sparse fully connected layers, but most of them target pre-trained, statically sparse weights. By contrast, PADE targets attention dynamic sparsity, which requires on-the-fly prediction, making zero-based sparsity methods ineffective. While some recent works explore activation sparsity [jang2021sparsity] or combine weight and activation sparsity [wu2023highlight, huang2023rm, wang2021dual], they still rely on zero-based sparsity, failing to address the argmax sparsity, which is the PADE targets.
Bit-serial computing accelerators. Prior works [wang2025mcbp, im2023sibia, han2024bitnn, yang2021fusekna, judd2016stripes, albericio2017bit, delmas2019bit, lo2023bit, sharify2019laconic, lee2018unpu, delmas2017dynamic, gondimalla2019sparten, li2022ristretto, parashar2017scnn, yang2023isosceles, lu2021distilling, kam2025panacea, li2025lut] accelerates NNs by exploiting bit-level sparsity [albericio2017bit, delmas2019bit, lu2021distilling, sharify2019laconic, yang2021fusekna] or dynamically reducing bit-width [gondimalla2019sparten, li2022ristretto, parashar2017scnn, yang2023isosceles]. However, these techniques are difficult to apply directly to sparse attention, as they typically rely on offline weight preprocessing, while sparse attention demands real-time, token-wise prediction. Further, the inherently low utilization of bit-level prediction further limits their effectiveness in attention acceleration. In contrast, PADE adopts a lightweight runtime pruning strategy, BUI-GF, and maximizes hardware efficiency via bidirectional bit sparsity and streamlined out-of-order execution.
VIII Conclusion
We propose PADE, a software-hardware co-design for dynamic sparse attention without relying on traditional sparsity predictors. Through bit-level early termination, dual-sided bit sparsity, dedicated bit-wise out-of-order execution, and an optimized tiling dataflow, PADE achieves up to $31.1\times$ energy efficiency over H100 and $5.1\times$ over SOTA accelerator Sanger.
Acknowledgments
This work was supported in part by the National Science and Technology Major Project under Grant 2022ZD0115200; in part by the NSFC under Grant 62125403, Grant U24A20234, Grant 92464302 and Grant U24B20164; in part by the Beijing S&T Project Z251100008425010; in part by Shanghai Municipal Science and Technology Major Project; the Natural Science Foundation of Jiangsu Province Basic Research Program under Grant BK20243042; in part by the Beijing National Research Center for Information Science and Technology; in part by the Northern IC Technology Innovation Center (Beijing) Co., Ltd under Grant QYJS20232801B; and in part by the Beijing Advanced Innovation Center for Integrated Circuits.

### v2

@@
 I Introduction
-Transformer models have achieved significant success in various fields, spanning from content generation [song2018situ, brown2020language, hendrycks2020measuring, lin2023videodirectorgpt] to computer vision [dosovitskiy2020image, you2023vitcod, bai2024seed].
+Transformer models have achieved significant success in various fields, spanning from content generation [107, 5, 47, 75] to computer vision [27, 145, 7].
 However, the self-attention mechanism used in Transformers suffers from quadratic time and memory complexity, limiting their scalability to long sequences.
@@
 Existing sparse attention works can be divided into two routes: static sparsity (SS) and dynamic sparsity (DS).
-SS [parmar2018image, child2019generating, qiu2019blockwise, zaheer2020big, beltagy2020longformer, jiang2024minference, dong2022cswin, katharopoulos2020transformers] relies on predefined sparse patterns that remain fixed during inference, lacking flexibility and often resulting in significant accuracy degradation [li2024large, fuad2023survey, kachris2025survey].
-In contrast, DS [correia2019adaptively, kitaev2020reformer, you2023vitcod, wang2023cta, song2024tsacc, cho2024sparc, park2024token, fan2022adaptable, dass2023vitality, li2020ftrans, wang2024sofa, qin2023fact, ham20203, ham2021elsa, lu2021sanger, qu2022dota, yang2022dtatrans, zhou2022energon, wang2025bitstopper, wang2021spatten, liu2022dynamic, wang2025lapa, li2022accelerating, wang2025mcbp] adapts the sparsity pattern at runtime, offering improved accuracy and flexibility, making it more suitable for a broader range of tasks and input types.
+SS [89, 16, 94, 146, 9, 56, 26, 61] relies on predefined sparse patterns that remain fixed during inference, lacking flexibility and often resulting in significant accuracy degradation [72, 34, 58].
+In contrast, DS [19, 63, 145, 117, 108, 17, 88, 29, 20, 69, 119, 93, 41, 42, 81, 95, 142, 150, 121, 116, 76, 122, 74, 125] adapts the sparsity pattern at runtime, offering improved accuracy and flexibility, making it more suitable for a broader range of tasks and input types.
 Figure 1: Comparison of (a) current DS works and (b) PADE.
@@
 Fig.
-2 (a) shows the power breakdown of dense attention and two representative DS accelerators: Sanger [lu2021sanger], SOFA [wang2024sofa], with varying executor bit-widths.
+2 (a) shows the power breakdown of dense attention and two representative DS accelerators: Sanger [81], SOFA [119], with varying executor bit-widths.
 Power consumption is categorized into executor and predictor components.
@@
 (b) Power ratio of predictor and executor versus SL with under 8-bit quantized executor.
-Takeaway: With the rapid advancement of Transformer quantization techniques, like GPTQ [frantar2022gptq], LLM.int8() [dettmers2022gpt3], SmoothQuant [xiao2023smoothquant] and Atom [zhao2024atom], there is a growing trend of adopting low-bit quantization in attention mechanisms.
+Takeaway: With the rapid advancement of Transformer quantization techniques, like GPTQ [33], LLM.int8() [25], SmoothQuant [137] and Atom [149], there is a growing trend of adopting low-bit quantization in attention mechanisms.
 In these cases, the predictor’s overhead increasingly offsets the benefits of sparsity.
@@
 Insights: The root cause of the excessive prediction cost stems from the decoupling between existing sparsity predictors and executors, which hinders the computational and memory access efforts paid in the predictor from being reused by the executor.
-To address this, we draw inspiration from bit-serial computing [judd2016stripes, albericio2017bit, lee2018unpu, chen2024bbs, kam2025panacea, im2023sibia, guo2025transitive], which separates an INT operation into multi-round bit-level steps.
+To address this, we draw inspiration from bit-serial computing [57, 1, 66, 15, 59, 52, 40], which separates an INT operation into multi-round bit-level steps.
 This motivates a unified design that integrates prediction and execution into a single computation stage, thereby eliminating the separate predictor and improving overall efficiency.
@@
 4(a), traditional DS works adopt a stage-splitting paradigm that decouples prediction from execution.
-Taking Sanger [lu2021sanger] as an example, the predictor takes the full 4-bit Key tensor to identify iQKs, while the executor separately fetches the corresponding KVs for precise computation.
+Taking Sanger [81] as an example, the predictor takes the full 4-bit Key tensor to identify iQKs, while the executor separately fetches the corresponding KVs for precise computation.
 However, this paradigm introduces two inefficiency sources: (1) For Keys related to iEQKs (e.g., K0), a single bit may suffice to identify their insignificance.
@@
 Fig.
-4(c) profiles the memory access and computation reduction achieved by different strategies in the attention modules across four layers from LLaMA-2 7B [touvron2023llama2].
+4(c) profiles the memory access and computation reduction achieved by different strategies in the attention modules across four layers from LLaMA-2 7B [113].
 On average, BSF can achieve $4.6\times$ higher memory access and $2.1\times$ more computation reduction, compared to traditional stage-splitting DS approaches, highlighting its significant energy efficiency advantage over current DS methods.
@@
 5 (b).
-Key idea We introduce the concept of a bit uncertainty interval (BUI), which defines the maximum fluctuation range of dot products across diverse bit-planes, via simple yet efficient bit-flipping.
+Key idea.
+Inspired the conservative margin concept [74, 88], we propose bit uncertainty interval-enabled guarded filtering (BUI-GF), a lightweight mechanism for simple yet accurate max-based pruning.
+BUI-GF characterizes the maximum possible fluctuation of dot products across bit planes using simple yet efficient bit flipping, enabling conservative yet hardware-efficient pruning decisions with minimal overhead.
 To support this, we design a scoreboard-based result-reuse PE to reduce BUI hardware overhead and redundant bit-plane memory accesses.
@@
 If true, the subsequent bit-planes for that Key are skipped; otherwise, the next bit-plane is required.
-However, due to DRAM’s dynamic precharge mechanism, loading each bit-plane typically incurs several dozen cycles [chang2016understanding, JEDEC_JESD235D].
+However, due to DRAM’s dynamic precharge mechanism, loading each bit-plane typically incurs several dozen cycles [12, 55].
 Naively stalling computation during data loading results in underutilized computational resources.
 Key idea.
-We propose bidirectional sparsity (BS), treating bit ‘1’ as a form of sparsity, thus ensuring load imbalance is below $50\%$.
-Building upon this, we introduce bit-wise out-of-order (OOE) execution, allowing the PE to process other bit-planes while avoiding memory access stalls.
-To support this, we propose a lightweight ANDer tree to reduce multiplexer’s overhead from BS sparsity, while the scoreboard-based PE facilitates partial sum buffering and reuse.
+Inspired by bidirectional sparsity [15] and XNOR-BNN-based formulations [35], we propose a two-pronged approach to mitigate resource underutilization.
+First, we introduce bidirectional, runtime-adaptive sparsity orchestration for the K matrix, which dynamically interprets bit ‘1’ as sparsity in coordination with queries, ensuring load imbalance remains below $50\%$.
+Building upon this, we further introduce bit-wise out-of-order (OOE) execution, allowing the PE to process other bit-planes while avoiding memory access stalls.
+To support this, we propose a temporal-reuse-based sparsity scheduler to alleviate runtime scheduling overhead, a lightweight ANDer tree to BS-induced multiplexing overhead, and a scoreboard-based PE that facilitates partial-sum buffering and reuse.
 (Challenge 3) IO inefficiency resulting from the conflict between tiling with the row-dependent pruning strategy.
@@
 A coarse solution is to enlarge on-chip SRAM, but this incurs significant area inefficiency.
-For example, with (PP=512, SS=2048), 5MB of SRAM is required, resulting in a 5.47 mm2 footprint under TSMC 28nm technology, which is $7.4\times$ and $8.9\times$ larger than the total area of SpAtten [wang2021spatten] and ELSA [ham2021elsa], respectively.
+For example, with (PP=512, SS=2048), 5MB of SRAM is required, resulting in a 5.47 mm2 footprint under TSMC 28nm technology, which is $7.4\times$ and $8.9\times$ larger than the total area of SpAtten [116] and ELSA [42], respectively.
 TABLE I: Summary for SOTA Attention Accelerators.
@@
 Computation Memory Free Support Level
-ELSA [ham2021elsa] $\times$ $\times$ $\times$ Value
-Sanger [lu2021sanger] $\times$ $\times$ $\times$ Value
-DOTA [qu2022dota] $\times$ $\times$ $\times$ Value
-DTATrans[yang2022dtatrans] Low ✓* $\times$ Value
-SpAtten [wang2021spatten] Low ✓* $\times$ Multi-bit
-Energon [zhou2022energon] $\times$ $\times$ $\times$ Multi-bit
-FACT [qin2023fact] $\times$ $\times$ $\times$ Value
-SOFA [wang2024sofa] Low $\times$ Value
+ELSA [42] $\times$ $\times$ $\times$ Value
+Sanger [81] $\times$ $\times$ $\times$ Value
+DOTA [95] $\times$ $\times$ $\times$ Value
+DTATrans[142] Low ✓* $\times$ Value
+SpAtten [116] Low ✓* $\times$ Multi-bit
+Energon [150] $\times$ $\times$ $\times$ Multi-bit
+FACT [93] $\times$ $\times$ $\times$ Value
+SOFA [119] Low $\times$ Value
 PADE Bit
@@
 Table I summarizes their features.
-The majority of existing works [ham20203, ham2021elsa, lu2021sanger, qu2022dota, yang2022dtatrans] focus on accelerating attention by alleviating computation overhead.
-For example, ELSA [ham2021elsa], Sanger [lu2021sanger], DOTA [qu2022dota], FACT [qin2023fact] adopt techniques like binary hashing, half-precision MSB, low-rank approximation, log-domain shifting to accelerate computation.
+The majority of existing works [41, 42, 81, 95, 142] focus on accelerating attention by alleviating computation overhead.
+For example, ELSA [42], Sanger [81], DOTA [95], FACT [93] adopt techniques like binary hashing, half-precision MSB, low-rank approximation, log-domain shifting to accelerate computation.
 However, these methods overlook memory optimization.
-While SpAtten [wang2021spatten].
-SOFA [wang2024sofa] realizes this challenge, their coarse-grained strategies, like hybrid quantization, and cross-stage tiling, fail to exploit fine-grained, bit-level optimizations.
+While SpAtten [116].
+SOFA [119] realizes this challenge, their coarse-grained strategies, like hybrid quantization, and cross-stage tiling, fail to exploit fine-grained, bit-level optimizations.
 Further, all current works rely on extra sparsity predictors, incurring substantial overhead.
-Notably, DTATrans [yang2022dtatrans] and SpAtten [wang2021spatten] guide sparsity using attention scores from the previous layer.
+Notably, DTATrans [142] and SpAtten [116] guide sparsity using attention scores from the previous layer.
 While this strategy partially reduces predictor overhead, it necessitates resource-intensive retrain to recover accuracy.
@@
 Second, bit-wise out-of-order execution (OOE) hides DRAM access latency.
-The core idea of BS is that 1 is also a form of sparsity. Without loss of generality, the dot-product between an $N$-element Query and Key can be formulated as:
+We extend the BS from static weight scenarios [15] to dynamic attention workloads.
+The core idea of BS is that bit value ‘1’ is also a form of sparsity.
+Without loss of generality, the dot-product between an $N$-element Query and Key can be formulated as:
 $\sum\nolimits_{j=0}^{N-1}q_{j}k_{j}=\sum\nolimits_{b=0}^{p-1}2^{b}\times\sum\nolimits_{j=0}^{N-1}q_{j}\times k_{j}^{b},$ (5)
@@
 This transformation ensures that at most $50\%$ of the bits are involved in computation, thus effectively improving the load balance across PEs.
+Notably, unlike prior work [15] that targets static weight compression or bit skipping, we extend bidirectional sparsity as a runtime load-balancing mechanism to bound PE imbalance in bit-serial QK execution, where a lightweight sparsity-aware scheduler (see §V-D) is crucial due to the highly dynamic, runtime-determined nature of attention workloads.
 […]
@@
 $softmax(x_{i})=\frac{\exp(x_{i})}{\sum_{j=0}^{N-1}\exp(x_{j})}\leq\frac{\exp(x_{i})}{\sum_{j\in subset}\exp(x_{j})}$ (7)
+Figure 10: (a) Illustration for head-tail interleaving updating. (b) Heavy updating cost across tiles. (c) Detailed procedure of the ISTA algorithm.
 To identify retained keys within a subset, a key is retained if all its bit planes are processed and it remains unpruned.
@@
 If a Key (i.e., $K_{0}$) remains unpruned after processing its least significant bit (LSB) plane, it is deemed an important token, and its index is stored in the Retained Key Board.
+Figure 11: (a) Overview of the PADE architecture. (b) Bit-wise PE lane. (c) BUI Generator. (d) BUI-GF Module. (e) Decision Unit.
 Driven by the tile-level sparsity decisions, the ISTA algorithm is detailed in Fig.
@@
 Specifically, each time the maximum is updated, it triggers one subtraction, one exponentiation and two multiplications of scalar and vector (see highlights in lines 11-12).
-We employ the arithmetic complexity model [brent2010modern, wang2021efficientMIMO, wang2023efficient] to normalize the complexity for different operations.
+We employ the arithmetic complexity model [10, 120, 118] to normalize the complexity for different operations.
 As profiled in Fig.
@@
 Moreover, the overhead becomes more as $B_{c}$ decreases, due to more frequent updates.
-Figure 10: (a) Illustration for head-tail interleaving updating. (b) Heavy updating cost across tiles. (c) Detailed procedure of the ISTA algorithm.
-Figure 11: (a) Overview of the PADE architecture. (b) Bit-wise PE lane. (c) BUI Generator. (d) BUI-GF Module. (e) Decision Unit.
 To minimize the overhead associated with the ‘max updating’, it is advantageous to prioritize dominant tokens.
 Note that the reduction in operations refers to the redundant computations shown in lines 11–12, rather than the comparison operations in line 9.
-To this end, we propose a head-tail interleaved updating strategy that exploits the locality property of attention: Recently generated tokens and the initial token typically exhibit higher weights than others [tian2023scan, jiang2024minference].
+To this end, we propose a head-tail interleaved updating strategy that exploits the locality property of attention: Recently generated tokens and the initial token typically exhibit higher weights than others [112, 56].
 As illustrated in Fig.
@@
 V-D Grouped Lightweight Sparsity ANDer Tree (GSAT)
+Figure 12: Temporal reuse architecture of bidirectional sparsity (BS) scheduler.
 To exploit bit sparsity, a direct design is to select and accumulate activations corresponding to non-zero bits.
@@
 17 (a).
-BS Scheduler.
-PADE adopts a low-cost BS scheduler to manage the operations within each PE, as depicted in Fig.
+Lightweight BS Scheduler.
+Built upon [15], PADE adopts a low-cost BS scheduler by reusing a critical priority encoder across time steps to orchestrate operations within each PE, as depicted in Fig.
 12.
+Unlike the straightforward design in [15], which instantiates multiple priority encoders in parallel, PADE temporally multiplexes a single priority encoder across successive time steps for index selection.
+This temporal reuse is enabled by PADE’s distinctive pipelined microarchitecture, where the QK-PU and V-PU operate in a staggered pipeline, effectively hiding the additional latency introduced by encoder reuse.
+The temporal reuse scheme helps PADE reduce 75% priority encoder overhead.
 To control the bit-serial dot product, the scheduler first identifies whether a bit plane of the Key vector contains more zeros or ones.
@@
 If such a bit exists, it is masked, and the remaining bits are propagated to the next time step.
-Otherwise, if the vector contains only zeros, the encoder asserts val=0 to disable the corresponding bit-serial multiplier in the PE, as depicted in Fig.
+Otherwise, if the vector contains only zeros, the encoder asserts Vx=0 to disable the corresponding bit-serial multiplier in the PE, as depicted in Fig.
 11 (b).
-Figure 12: Detailed architecture of bidirectional sparsity (BS) scheduler.
+Figure 13: Reuse-aware reorder scheduler (RARS).
+TABLE II: Accuracy of Different Transformer Models with MXINT8, FP16, INT8 and PADE Configurations (S: Standard, A: Aggressive).00footnotetext:
+Model LlaMa2-7B LlaMa3-8B OPT1B3 Bloom1B7 Qwen7B ViT-L/16 PVT
+Task Dolly Wikili. MBPP Wiki2 MMLU Winog. Dolly Wikili. MBPP Wiki2 MMLU Winog. Wikili. MBPP Wikili. MBPP Wikili. MBPP Image VTAB Image VTAB
+MXINT8 36.5 39.3 17.5% 5.63 35.2% 69.8% 40.9 43.6 23.3% 5.01 42.2% 75.1% 36.1 11.9% 44.6 16.3% 46.8 30.5% 85.5% 72.8% 89.7% 77.5%
+FP16 36.4 39.1 17.5% 5.71 35.1% 69.4% 40.8 42.7 21.8% 5.11 41.2% 74.2% 36.2 11.9% 44.3 16% 46.6 30% 85.3% 72.7% 89.4% 77.3%
+INT8 36.4 38.9 17.2% 5.73 34.7% 69.3% 40.7 42.7 21.6% 5.13 40.9% 73.7% 35.9 11.6% 44.1 15.7% 46.4 29.2% 85.3% 72.5% 89.3% 77.1%
+PADE (S) 36.3 38.9 17.2% 5.75 34.6% 69.2% 40.6 42.6 21.5% 5.13 40.7% 73.7% 35.9 11.5% 44.0 15.6% 46.3 29.2% 85.3% 72.5% 89.3% 77.1%
+PADE (A) 36.1 38.4 16.5% 5.80 34.1% 68.7% 40.5 42.0 21.0% 5.19 40.2% 72.8% 35.3 11.0% 43.6 15.2% 45.9 28.4% 84.9% 72.4% 89.1% 76.8%
+MMLU, WinoGrande, MBPP, Imagenet, VTAB are evaluated by accuracy. Dolly, Wikilingua are evaluated by ROUGE-1. Wikitext2 is evaluated by PPL, where lower is better.
+Figure 14: Normalized computation and memory access across diverse Transformer models and tasks. Spatten* performs additional fine-tuning.
 V-E Reuse-Aware Reorder Scheduling (RARS)
@@
 VI-A Experimental Setup
-Baseline comparisons: We compare PADE with five SOTA attention accelerators: Sanger [lu2021sanger], Spatten [wang2021spatten], Energon [zhou2022energon], DOTA [qu2022dota], SOFA [wang2024sofa].
-For fair comparison, all designs are normalized to a 28nm process and evaluated under identical conditions: PE arrays occupy the same area as PADE and work in 800 MHz, on-chip SRAM is set to 352kB, and peak HBM bandwidth is 256 GB/s, with 4 pj/bit [o2017fine].
-Figure 13: Reuse-aware reorder scheduler (RARS).
-TABLE II: Accuracy of Different Transformer Models with MXINT8, FP16, INT8 and PADE Configurations (S: Standard, A: Aggressive).00footnotetext:
-Model LlaMa2-7B LlaMa3-8B OPT1B3 Bloom1B7 Qwen7B ViT-L/16 PVT
-Task Dolly Wikili. MBPP Wiki2 MMLU Winog. Dolly Wikili. MBPP Wiki2 MMLU Winog. Wikili. MBPP Wikili. MBPP Wikili. MBPP Image VTAB Image VTAB
-MXINT8 36.5 39.3 17.5% 5.63 35.2% 69.8% 40.9 43.6 23.3% 5.01 42.2% 75.1% 36.1 11.9% 44.6 16.3% 46.8 30.5% 85.5% 72.8% 89.7% 77.5%
-FP16 36.4 39.1 17.5% 5.71 35.1% 69.4% 40.8 42.7 21.8% 5.11 41.2% 74.2% 36.2 11.9% 44.3 16% 46.6 30% 85.3% 72.7% 89.4% 77.3%
-INT8 36.4 38.9 17.2% 5.73 34.7% 69.3% 40.7 42.7 21.6% 5.13 40.9% 73.7% 35.9 11.6% 44.1 15.7% 46.4 29.2% 85.3% 72.5% 89.3% 77.1%
-PADE (S) 36.3 38.9 17.2% 5.75 34.6% 69.2% 40.6 42.6 21.5% 5.13 40.7% 73.7% 35.9 11.5% 44.0 15.6% 46.3 29.2% 85.3% 72.5% 89.3% 77.1%
-PADE (A) 36.1 38.4 16.5% 5.80 34.1% 68.7% 40.5 42.0 21.0% 5.19 40.2% 72.8% 35.3 11.0% 43.6 15.2% 45.9 28.4% 84.9% 72.4% 89.1% 76.8%
-MMLU, WinoGrande, MBPP, Imagenet, VTAB are evaluated by accuracy. Dolly, Wikilingua are evaluated by ROUGE-1. Wikitext2 is evaluated by PPL, where lower is better.
-Figure 14: Normalized computation and memory access across diverse Transformer models and tasks. Spatten* performs additional fine-tuning.
-Benchmarks: We evaluate PADE on several representative Transformer models across NLP and CV tasks.
-For NLP tasks, Llama2-7B [touvron2023llama2], Llama3-8B [grattafiori2024llama], Qwen7B [bai2023qwen], Bloom1B7 [le2022bloom] and OPT1B3 [zhang2022opt], for six tasks.These tasks include language modeling (Wikitext-2 (S=2k) [merity2016pointer], Wikilingua (S=2k) [ladhak-wiki-2020], Winogrande (S=0.25k) [sakaguchi2021winogrande]), language understanding (MMLU, S=0.5k) [hendrycks2020measuring], code generation MBPP (S=1k) [austin2021program], long context processing dolly (S=15k) [conover2023free].
-For CV tasks, we choose ViT-L/16 (S=576) [dosovitskiy2020image] and PVT (S=3k) [wang2021pyramid] on ImageNet-1k [deng2009imagenet] and JFT [sun2017revisiting] classification.
-Quantization Accuracy.
-All pre-trained models are sourced from Pytorch [paszke2017automatic] and HuggingFace [wolf2020transformers].
-INT8 baselines derived via post-training quantization, where only the weights and activations (QKVs) are quantized to INT8, while non-linear operators (e.g., softmax) remain in FP16 precision.
-As shown in Table.
-II, the INT8 baseline incurs less than a $1\%$ average accuracy drop from FP16, thus confirming its validity.
-Hardware Evaluation: Table III lists the hardware configuration of PADE.
-We implement the RTL design for PADE and utilize Synopsys DC on TSMC 28nm CMOS technology to estimate the logic area and power.
-The power, area, and read/write bandwidth of on-chip SRAM buffers are estimated through CACTI [muralimanohar2009cacti].
-Off-chip HBM modeling involves simulating access patterns and row activation under various data layouts (Fig.
-22), capturing HBM’s burst behavior.
-We derive memory latency from Ramulator [kim2015ramulator], and estimate IO power following the methodology in [cavigelli2016origami, andri2016yodann, wang2017energy].
-We extract each stage’s cycles by simulating the RTL with Verilator [snyder2004verilator], and use a custom cycle-level simulator to evaluate performance.
+Baseline comparisons: We compare PADE with five SOTA attention accelerators: Sanger [81], Spatten [116], Energon [150], DOTA [95], SOFA [119].
+For fair comparison, all designs are normalized to a 28nm process and evaluated under identical conditions: PE arrays occupy the same area as PADE and work in 800 MHz, on-chip SRAM is set to 352kB, and peak HBM bandwidth is 256 GB/s, with 4 pj/bit [86].
 TABLE III: On-chip Hardware and Off-chip DRAM Configurations of PADE
@@
 […]
-GPU comparison We benchmark on an Nvidia H100 using SOTA TensorRT-LLM [Nvidia2023] with FlashAttention3 [shah2024flashattention].
+Benchmarks: We evaluate PADE on several representative Transformer models across NLP and CV tasks.
+For NLP tasks, Llama2-7B [113], Llama3-8B [38], Qwen7B [6], Bloom1B7 [65] and OPT1B3 [147], for six tasks.These tasks include language modeling (Wikitext-2 (S=2k) [83], Wikilingua (S=2k) [28], Winogrande (S=0.25k) [99]), language understanding (MMLU, S=0.5k) [47], code generation MBPP (S=1k) [4], long context processing dolly (S=15k) [18].
+For CV tasks, we choose ViT-L/16 (S=576) [27] and PVT (S=3k) [130] on ImageNet-1k [24] and JFT [110] classification.
+Quantization Accuracy.
+All pre-trained models are sourced from Pytorch [90] and HuggingFace [134].
+INT8 baselines derived via post-training quantization, where only the weights and activations (QKVs) are quantized to INT8, while non-linear operators (e.g., softmax) remain in FP16 precision.
+As shown in Table.
+II, the INT8 baseline incurs less than a $1\%$ average accuracy drop from FP16, thus confirming its validity.
+Hardware Evaluation: Table III lists the hardware configuration of PADE.
+We implement the RTL design for PADE and utilize Synopsys DC on TSMC 28nm CMOS technology to estimate the logic area and power.
+The power, area, and read/write bandwidth of on-chip SRAM buffers are estimated through CACTI [84].
+Off-chip HBM modeling involves simulating access patterns and row activation under various data layouts (Fig.
+22), capturing HBM’s burst behavior.
+We derive memory latency from Ramulator [62], and estimate IO power following the methodology in [11, 2, 133].
+We extract each stage’s cycles by simulating the RTL with Verilator [106], and use a custom cycle-level simulator to evaluate performance.
+Figure 15: (a)(b) Accuracy comparison with current sparse attention methods. (c) Speedup and efficiency gain comparison.
+GPU comparison We benchmark on an Nvidia H100 using SOTA TensorRT-LLM [85] with FlashAttention3 [100].
 To exclude the software overhead, we measure execution time with cudaEvent, isolating GPU execution from CPU interference.
@@
 For batch size, we will select the configuration from [8, 128] that maximizes GPU computational efficiency based on the sequence length of each dataset.
-Figure 15: (a)(b) Accuracy comparison with current sparse attention methods. (c) Speedup and efficiency gain comparison.
+Figure 16: (a) Latency reduction for BUI-GF, BS-OOE and ISTA. (b) Exploring the trade-off between accuracy and sparsity.
 VI-B Algorithm Performance
@@
 Fig.
-15 (a)(b) compares the accuracy of three representative software-only sparse attention methods, two predictor-free works (SpAtten [wang2021spatten], DTATrans [yang2022dtatrans]), and PADE.
+15 (a)(b) compares the accuracy of three representative software-only sparse attention methods, two predictor-free works (SpAtten [116], DTATrans [142]), and PADE.
 The Sparsity Level denotes the ratio between the computation cost of sparse execution (prediction + actual computation) and that of dense execution.
-StreamingLLM [xiao2023efficient] adopts a static sparsity pattern—retaining only the initial and recent tokens, while MInference [jiang2024minference], and DoubleSparsity [yang2024post] rely on runtime sparsity prediction.
+StreamingLLM [138] adopts a static sparsity pattern—retaining only the initial and recent tokens, while MInference [56], and DoubleSparsity [141] rely on runtime sparsity prediction.
 Four key observations are made: 1) StreamingLLM performs the worst, due to its lack of adaptivity in capturing input-dependent sparsity patterns.
@@
 […]
-Figure 16: (a) Latency reduction for BUI-GF, BS-OOE and ISTA. (b) Exploring the trade-off between accuracy and sparsity.
+Figure 17: Exploration of optimal (a) Sub-group size. (b) Scoreboard entry.
 Computation Reduction.
@@
 The key architectural parameters in PADE are decided by detailed workload profiling and DSE exploration.
-Figure 17: Exploration of optimal (a) Sub-group size. (b) Scoreboard entry.
 (1) Computational throughput: The INT-8 throughput ratio between QK-PU and V-PU is set to 8:1 (see Table III), derived from the typical QK-to-SV computation ratio observed in LLM workloads due to sparsity.
@@
 23 (a) presents a detailed breakdown of execution time across varying numbers of PE lanes to illustrate load balance effects.
-To better show PADE’s advantages, we compare it against a SOTA bit accelerator, BitWave [shi2024bitwave], which leverages bit-flipping to enhance bit-plane sparsity.
+To better show PADE’s advantages, we compare it against a SOTA bit accelerator, BitWave [104], which leverages bit-flipping to enhance bit-plane sparsity.
 Since each PE lane integrates multiple bit-serial multipliers, intra-PE stall arises when certain multipliers handle more effective bits.
@@
 Extension for MXINT.
-The micro-scale format performs fine-grained quantization along the channel dimension by grouping data into 32-element segments [rouhani2023microscaling].
+The micro-scale format performs fine-grained quantization along the channel dimension by grouping data into 32-element segments [96].
 PADE ensures compatibility by applying group-wise scaling to the bit uncertain interval (BUI) (§IV-A).
@@
 The resulting BUI is then expanded using the quantization factor, eg., $\Delta_{\rm Q1}\Delta_{\rm K1}/\Delta_{\rm A}$.
-❷ Finally, the max and min BUI values across all groups are aggregated to compute the overall BUI.
+Finally, the max and min BUI values across all groups are aggregated to compute the overall BUI.
 This method can be extended to dot products of arbitrary length.
 Extension for FP formats.
-Prior studies [sheng2023flexgen, xiao2023smoothquant, zhao2024atom] have demonstrated that the K and V tensors are highly amenable to quantization because the softmax normalization in self-attention naturally suppresses quantization noise.
+Prior studies [103, 137, 149] have demonstrated that the K and V tensors are highly amenable to quantization because the softmax normalization in self-attention naturally suppresses quantization noise.
 This property enables safe quantization to INT8 or even INT4 with negligible accuracy degradation.
-Motivated by this, when queries operate in FP format, PADE converts the INT-FP computation into a bit-serial form through exponent alignment, following methodologies adopted in prior works [chen2025bitmod, jang2024figna, fang2025anda].
+Motivated by this, when queries operate in FP format, PADE converts the INT-FP computation into a bit-serial form through exponent alignment, following methodologies adopted in prior works [14, 53, 31].
 The resulting bit-serial execution is fully compatible with PADE’s existing processing mechanism.
@@
 While PADE significantly advances sparse-attention acceleration by eliminating the sparsity predictor through unified bit-serial stage fusion, several limitations remain for future works: (1) As the sequence length of large language models (LLMs) continues to scale, the need for distributed attention becomes inevitable.
-Extending PADE to distributed scenarios, especially in emerging wafer-scale architectures [wang2026WATOS, wang2026TEMP, wang2024tmac, he2025waferllm, hu2024wafer, tang2025moentwine], remains an open problem.
+Extending PADE to distributed scenarios, especially in emerging wafer-scale architectures [124, 126, 128, 45, 50, 111], remains an open problem.
 (2) Beyond single-bit granularity, exploring multi-bit stage fusion may provide a more favorable efficiency–accuracy trade-off, which represents a promising direction for future PADE enhancements.
@@
 Efficient Attention Accelerators.
-Numerous attention accelerators [wang2025bitstopper, wang2025lapa, wang2025beta, li2020ftrans, ham20203, ham2021elsa, qu2022dota, fan2023taskfusion, yang2022dtatrans, hong2022dfx, liu2024hsconn, wang2024sofa, wang2022energy, fang2022algorithm, yazdanbakhsh2022sparse, li2022accelerating, lu2021sanger, wang2021spatten, zhou2022energon, qin2023fact, shen2022salo, fan2022adaptable, bai2024swat, zhao2024hardware, tuli2023acceltran, you2023vitcod, shen2022salo] have been proposed.
-Early works, such as A3 [ham20203] and ELSA [ham2021elsa] accelerate computation via approximation techniques.
+Numerous attention accelerators [121, 122, 123, 69, 41, 42, 95, 30, 142, 49, 77, 119, 131, 32, 144, 74, 81, 116, 150, 93, 102, 29, 8, 148, 114, 145, 102] have been proposed.
+Early works, such as A3 [41] and ELSA [42] accelerate computation via approximation techniques.
 Recent efforts have shifted to jointly optimizing computation and memory.
-Energon [zhou2022energon] and SOFA [wang2024sofa] adopt fine-grained filtering and FlashAttention-optimized tiling techniques to alleviate memory overhead.
+Energon [150] and SOFA [119] adopt fine-grained filtering and FlashAttention-optimized tiling techniques to alleviate memory overhead.
 However, they still rely on extra sparsity predictors, which become de facto latency and power bottlenecks after sparsification.
@@
 Neural network (NN) accelerator with sparsity.
-Numerous accelerators [wang2021efficient, hojabr2021spaghetti, wang2018low, gondimalla2019sparten, asgari2020alrescha, deng2021gospa, qin2020sigma, gudaparthi2022candles, fang2022algorithm, hanson2022cascading, lew2022anticipating, li2022ristretto, li2021escalate, li2022accelerating, liu2022s2ta, pavon2021via, rucker2021capstan, sadi2019efficient, walia2021fast, yang2020procrustes, Gon2023Eureka, wu2022sparseloop, wu2023highlight, shin2022griffin] exploit sparsity to accelerate NN inference.
-General-purpose sparse tensor accelerators [srivastava2020matraptor, mahmoud2020tensordash, kwon2019tensordimm, hegde2019extensor, kanellopoulos2019smash, chen2020tpspmv] support operations on sparse fully connected layers, but most of them target pre-trained, statically sparse weights.
+Numerous accelerators [127, 48, 129, 36, 3, 23, 92, 39, 32, 44, 68, 70, 73, 74, 78, 91, 97, 98, 115, 139, 37, 136, 135, 105] exploit sparsity to accelerate NN inference.
+General-purpose sparse tensor accelerators [109, 82, 64, 46, 60, 13] support operations on sparse fully connected layers, but most of them target pre-trained, statically sparse weights.
 By contrast, PADE targets attention dynamic sparsity, which requires on-the-fly prediction, making zero-based sparsity methods ineffective.
-While some recent works explore activation sparsity [jang2021sparsity] or combine weight and activation sparsity [wu2023highlight, huang2023rm, wang2021dual], they still rely on zero-based sparsity, failing to address the argmax sparsity, which is the PADE targets.
+While some recent works explore activation sparsity [54] or combine weight and activation sparsity [135, 51, 132], they still rely on zero-based sparsity, failing to address the argmax sparsity, which is the PADE targets.
 Bit-serial computing accelerators.
-Prior works [wang2025mcbp, im2023sibia, han2024bitnn, yang2021fusekna, judd2016stripes, albericio2017bit, delmas2019bit, lo2023bit, sharify2019laconic, lee2018unpu, delmas2017dynamic, gondimalla2019sparten, li2022ristretto, parashar2017scnn, yang2023isosceles, lu2021distilling, kam2025panacea, li2025lut] accelerates NNs by exploiting bit-level sparsity [albericio2017bit, delmas2019bit, lu2021distilling, sharify2019laconic, yang2021fusekna] or dynamically reducing bit-width [gondimalla2019sparten, li2022ristretto, parashar2017scnn, yang2023isosceles].
+Prior works [125, 52, 43, 140, 57, 1, 22, 79, 101, 66, 21, 36, 70, 87, 143, 80, 59, 71] accelerates NNs by exploiting bit-level sparsity [1, 22, 80, 101, 140] or dynamically reducing bit-width [36, 70, 87, 143].
 However, these techniques are difficult to apply directly to sparse attention, as they typically rely on offline weight preprocessing, while sparse attention demands real-time, token-wise prediction.